OpenStudy (anonymous):

The Utah Ski Club sells calendars to raise money. The profit P, in cents, from selling x calendars is given by the equation P(x)=360x-x2. A.) find how many calendars must be sold to maximize profit. B.) find the maximum profit.

OpenStudy (amistre64):

utah?

OpenStudy (amistre64):

the vertex of will answer all your questions :)

OpenStudy (amistre64):

2x -360 = 0 x = 360/2 = 180

OpenStudy (anonymous):

maximum is at vertex which we now know is \[\frac{-b}{2a}\]

OpenStudy (anonymous):

180 i think is the answer for the first one.

OpenStudy (anonymous):

in this case it is 180

OpenStudy (anonymous):

then replace x by 180 and get your answer

OpenStudy (anonymous):

the profit is 180^2

OpenStudy (amistre64):

which after taxes is abot -.32

OpenStudy (anonymous):

actually profit is \[P(180)=360\times 180-180^2\]

OpenStudy (anonymous):

32400 which according to amistre rounds to -.32

OpenStudy (amistre64):

i got me a new book from library 2day

OpenStudy (anonymous):

don't forget the licensing fee

OpenStudy (anonymous):

and the bribe

OpenStudy (anonymous):

wadja get?

OpenStudy (amistre64):

differential equations , second edition. its the sequal to differential equation, first edition lol